# Fair Summarization: Bridging Quality and Diversity in Extractive Summaries

### v3

@@ 5.3 Evaluation Metrics
  • SUPERT: SUPERT (Gao et al.,, 2020) evaluates the quality of a summary by measuring its semantic similarity with a pseudo reference summary. It employs contextualized embeddings and soft token alignment techniques, providing an in-depth analysis of the semantic fidelity of generated summaries.
  • BLANC: BLANC (Vasilyev et al.,, 2020) is a reference-less metric that measures the improvement in a pretrained language model’s performance during language understanding tasks when given access to a summary.
  • SummaQA: SummaQA (Scialom et al.,, 2019) employs a question-answering model based on BERT to answer cloze-style questions using the system-generated summaries, providing insights into the summarization’s factual accuracy and coherence.
  • BARTScore: BARTScore (Yuan et al.,, 2021) is a parameter- and data-efficient metric that supports the evaluation of generated text from multiple perspectives, including informativeness and coherence.
  • UniEval: UniEval (Zhong et al.,, 2022) is a unified multi-dimensional evaluator that reframes natural language generation evaluation as a Boolean Question Answering (QA) task, guiding the model with different questions to evaluate from multiple dimensions. It is reference-free in three dimensions (coherence, consistency, fluency), but not relevance. For our evaluation, we focused on the reference-free dimensions of UniEval and reported the overall average performance.
  • Fairness (F): To align fairness with the quality metrics, we define $F=1-\text{RG}$, where larger values represent better fairness. The Representation Gap (RG) metric (Olabisi et al.,, 2022) assesses the fairness of summaries by measuring the balance in the representation of different groups.
@@ 7.1 Limitations
First, our methods focus on extractive summarization, which, while preserving input fidelity, may not capture the semantic richness of abstractive methods (Lebanoff et al.,, 2019). Extending our approach to abstractive models presents additional challenges, particularly in balancing fairness with coherence and fluency.
@@
Third, our work focuses on monolingual inputs, specifically in English. Future research could extend these methods to multilingual inputs, where additional factors such as language diversity and cross-lingual transfer, as highlighted by Bagheri Nezhad and Agrawal, (2024) and Bagheri Nezhad et al., (2025), would need to be addressed to ensure fairness across languages.
Additionally, while we employ standard quality and fairness metrics, they do not fully capture subjective factors such as readability or user trust. Human evaluation could provide deeper insights into the practical implications of fairness and quality.
Finally, the computational complexity of fair clustering and large language models may limit scalability in real-time or resource-constrained environments.
Despite these challenges, our work marks a significant step toward fairer summarization models, and addressing these limitations could enhance the robustness of fairness in NLP.
Acknowledgments and Disclosure of Funding
We would like to thank Aravind Inbasekaran for his valuable assistance throughout this project. This research has been supported by the National Science Foundation under Grant No. AF 2311397 and CRII:RI 2246174.
@@ A.1 Fair Extract Formal Algorithmic Processes
A.2 Sample of Dataset
Table 4 presents a sample of the dataset used in this study, containing tweets from different social groups about Chicago. Each entry indicates the group (White, African-American (AA), or Hispanic(Hisp)) and the corresponding tweet.
Group Tweet
White Turns out not all White Castles are the same. Why do you push me away Chicago?!
AA "I mean I’m from Chicago. I’ll cheer for the Bears, but I’m a bigger 49ers fan."
White Nothing makes me happier than seeing the Bulls win _____ #ChicagoBasketball #Bullieve
White If you see on the news something about the Chicago Kitchen Clown Bandits, then it will be referring to me, my friend Eten, and I.
AA Truuu we tryna find sum to do too.. I dnt wanna b n Chicago if ain’t nobody here.
White Oh yeah.. I’m good. Hangin’ up here in Chicago today. :)
Hisp You girls have a safe flight.! See you in Chicago (:
… (Dataset continues with more examples)
A.3 Impact of Varying Fairness Weight on Composite Metrics

### v5

@@ 5.3 Evaluation Metrics
  • SUPERT: SUPERT (Gao et al., 2020) evaluates the quality of a summary by measuring its semantic similarity with a pseudo reference summary. It employs contextualized embeddings and soft token alignment techniques, providing an in-depth analysis of the semantic fidelity of generated summaries.
  • BLANC: BLANC (Vasilyev et al., 2020) is a reference-less metric that measures the improvement in a pretrained language model’s performance during language understanding tasks when given access to a summary.
  • SummaQA: SummaQA (Scialom et al., 2019) employs a question-answering model based on BERT to answer cloze-style questions using the system-generated summaries, providing insights into the summarization’s factual accuracy and coherence.
  • BARTScore: BARTScore (Yuan et al., 2021) is a parameter- and data-efficient metric that supports the evaluation of generated text from multiple perspectives, including informativeness and coherence.
  • UniEval: UniEval (Zhong et al., 2022) is a unified multi-dimensional evaluator that reframes natural language generation evaluation as a Boolean Question Answering (QA) task, guiding the model with different questions to evaluate from multiple dimensions. It is reference-free in three dimensions (coherence, consistency, fluency), but not relevance. For our evaluation, we focused on the reference-free dimensions of UniEval and reported the overall average performance.
  • Fairness (F): To align fairness with the quality metrics, we define $F=1-\text{RG}$, where larger values represent better fairness. The Representation Gap (RG) metric (Olabisi et al., 2022) assesses the fairness of summaries by measuring the balance in the representation of different groups. We define perfect fairness as $F=1$, meaning the summary includes an equal number of documents from each social group. This metric only captures numerical balance and does not address other dimensions such as content diversity or semantic nuances, which we leave for future work.
@@ 8 Limitations
First, our methods focus on extractive summarization, which, while preserving input fidelity, may not capture the semantic richness of abstractive methods (Lebanoff et al., 2019). Extending our approach to abstractive models presents additional challenges, particularly in balancing fairness with coherence and fluency.
@@
Third, our work focuses on monolingual inputs, specifically in English. Future research could extend these methods to multilingual inputs, where additional factors such as language diversity and cross-lingual transfer Bagheri Nezhad and Agrawal (2024); Bagheri Nezhad et al. (2025), would need to be addressed to ensure fairness across languages.
Additionally, while we employ standard quality and fairness metrics, they do not fully capture subjective factors such as readability or user trust. Human evaluation could provide deeper insights into the practical implications of fairness and quality. Also, our evaluation primarily relies on quantitative metrics, we acknowledge that a deeper qualitative error analysis—examining specific examples and error cases—would further illuminate the limitations of fairness-aware summarization, and we consider this an important direction for future investigation.
Finally, the computational complexity of fair clustering and large language models may limit scalability in real-time or resource-constrained environments
9 Acknowledgments
We would like to thank Aravind Inbasekaran for his valuable assistance throughout this project. We also appreciate the constructive feedback provided by the anonymous reviewers. This research has been supported by the National Science Foundation under Grants AF 2311397 and CRII:RI 2246174.
@@ A.1 Fair Extract Formal Algorithmic Processes
A.2 Impact of Varying Fairness Weight on Composite Metrics